Present ages of Sameer and Anand are in the ratio of 5 : 4 respectively. Three years hence, the ratio of their ages will become 11 : 9 respectively. What is Anand's present age in years?






















A. 24 B. 27
C. 40 D. Cannot be determined
E. None of these

Grade:11

1 Answers

Simran Bhatia
10 years ago

Answer: Option A

Explanation:

Let the present ages of Sameer and Anand be 5x years and 4x years respectively.

Then, 5x + 3 = 11
4x + 3 9

 9(5x + 3) = 11(4x + 3)

 45x + 27 = 44x + 33

 45x - 44x = 33 - 27

 x = 6.

 Anand''s present age = 4x = 24 years.